OpenStudy (anonymous):

Use the rule of 70 to approximate the growth rate when the doubling time is: a. 5740 years Round the answer to three decimal places. Growth rate: _______% per year b. 4 seconds Enter the exact answer. Growth rate: _____% per second

OpenStudy (anonymous):

The rule of 70 says that \[ \text{Doubling Time} \approx \frac{70}{\text{Growth Rate}} \] Or, for you, \[\text{Growth Rate} \approx \frac{70}{\text{Doubling Time}} \] Where the doubling time is measured in years.

OpenStudy (anonymous):

so would that be.... G.Rate = 70/5740 G.Rate then would equal =1.29%

OpenStudy (anonymous):

and G.Rate = 70/4=17.5?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@Mertsj , @Luis_Rivera ,@Zarkon ,@jim_thompson5910 May you help ?

OpenStudy (mertsj):

Did you change the 4 seconds to years?

OpenStudy (anonymous):

I do not know how to do that....

OpenStudy (anonymous):

Would I still use the formula that they first gave me? the Growth Rate = 70/doubling time?

OpenStudy (anonymous):

and for the first one it should be more easier right since its already in years? but we want the growth percentage though..

OpenStudy (anonymous):

how do i convert the 551, 880,000 into a %?

OpenStudy (mertsj):

I just reread the problem. Scratch that because it says they want the growth rate per second.

OpenStudy (anonymous):

oh okay thats fine...

OpenStudy (anonymous):

Mertsj?

OpenStudy (mertsj):